OpenStudy (anonymous):

solve the following equation by clearing fractions, show work: 7/3-5/2x = 8+2x

OpenStudy (jlastino):

7/3 - 5/2x = 8+2x Multiply both sides by 6x 6x(7/3 - 5/2x) = 8+2x(6x) 2x(7) - 3(5) = 48x +12x^2 14x-15=48x +12x^2 12x^2+34x+15 = 0 Solve for x using quad. equation

OpenStudy (anonymous):

thanks Jlastino you been very helpful

OpenStudy (jlastino):

glad to help :D